OpenStudy (anonymous):

Line AB is dilated by a scale factor of 3 to form line A' B' . Point O is the center of dilation, and point O lies on line AB. If the slope of line AB is 3, what can be said about line A'B'? A. The slope of line A'B' is 6, but A'B' does not pass through o. B.The slope of line A'B' is 9, and line A'B' passes through o. C. The slope of line A'B' is 9, but line A'B' does not pass through o. D. The slope of line A'B' is 3 and line A'B' passes through o. E. The slope of line A'B' is 3, but line A'B' does not pass through o.

jimthompson5910 (jim_thompson5910):



jimthompson5910 (jim_thompson5910):

What happens when you dilate (using center point O)?

OpenStudy (anonymous):

The object becomes larger.

jimthompson5910 (jim_thompson5910):

good, the scale factor is 3, so that means OA' = 3*OA OB' = 3*OB